Twice a number added to half of itself equals 20.Find the number
Answer by math_tutor2020(3816)


x = some unknown number
2x = twice the number
x/2 = 0.5x = half of the number
2x + 0.5x = adding the two previous expressions
2x + 0.5x = 20

Solving for x
2x + 0.5x = 20
(2 + 0.5)x = 20
2.5x = 20
x = 20/(2.5)
x = 8
The unknown number is 8

2x = 2*8 = 16 is the result after doubling
x/2 = 8/2 = 4 is the result after halving
Sum = 16+4 = 20 which confirms we have the correct x value.

Answer: 8